Dawson Toys, Ltd., produces a toy called the Maze. The company has recently established a standard cost system to help control costs and has established the following standards for the Maze toy:

Direct materials: 6 microns per toy at $0.32 per micron

Direct labor: 1.4 hours per toy at $6.60 per hour

During July, the company produced 4,500 Maze toys. The toy's production data for the month are as follows:

Direct materials: 77,000 microns were purchased at a cost of $0.28 per micron. 43,250 of these microns were still in inventory at the end of the month.

Direct labor: 6,700 direct labor-hours were worked at a cost of $48,910.

Required:

1. Compute the following variances for July: (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations. Round final answer to the nearest whole dollar amount.)

a. The materials price and quantity variances.

b. The labor rate and efficiency variances.

Material price variance: _____ F/U

Material quantity variance: _____F/U

Labor rate variance: _____F/U

Labor efficiency variance: ____ F/U
Answer #1

a. Materials price variance = ( Actual Price - Standard Price ) x Actual Quantity Used = $ ( 0.28 - 0.32) x 77,000 microns = $ 3,080 F

Materials quantity variance = ( Actual quantity used - Standard quantity allowed for actual output ) x Standard Price = ( 33,750 - 4,500 x 6 ) x $ 0.32 = $ 2,160 U

b. Labor rate variance = ( Actual labor hour rate - Standard labor hour rate ) x Actual hours used = $ ( 7.30 - 6.60) x 6,700 hours = $ 4,690 U

Labor efficiency variance = ( Actual hours used - Standard hours allowed for actual output ) x Standard labor hour rate = ( 6,700 - 4,500 x 1.40) x $ 6.60 = $ 2,640 U